# A Survey on Vision-Language-Action Models for Embodied AI

### v2

@@ III-B1 Early Non-Transformer Control Policies
BC-Z [130] processes two types of task instructions: a language instruction or a human demonstration video. A language instruction undergoes encoding using the USE language encoder [172], while the video is encoded with ResNet18 [29]. The environment is presented to the model in the form of an RGB image, also encoded using ResNet18. Then the instruction embedding and the image embedding are combined through the FiLM layer [131], culminating in the generation of actions. This conditional policy is asserted to exhibit zero-shot task generalization to unseen tasks.
@@
Universal Policy (UniPi) [158] treats the decision-making problem as a text-conditioned video generation problem. To predict actions, UniPi generates a video based on a given text instruction, and actions are extracted from the frames of the video through inverse dynamics. This innovative policy-as-video formulation offers several advantages, including enhanced generalization across diverse robot tasks and the potential for knowledge transfer from internet videos to real robots.
@@
Figure 4: Three common architectures of Transformer-based control policies are characterized by their vision-language fusion methods. FiLM layers (Hadamard product $\odot$) are employed to fuse language and vision in RT-1 models. Some action decoders utilize cross-attention to condition on the instruction. Concatenation ($\oplus$) is the prevailing method of vision-language fusion for action decoders.
III-B2 Transformer-based Control Policies
BeT [173], an imitation learning variant of the Trajectory Transformer, addresses the noisy and multi-modal nature of non-expert human demonstration data by using k-means-based action discretization coupled with continuous action correction. C-BeT [174] extends BeT by adding a target frame or demonstration as the goal specification. VQ-BeT [175] further improves BeT by incorporating vector quantization instead of k-means, enabling better modeling of long-range actions.
RoboCat [127] proposes a self-improvement process designed to enable an agent to rapidly adapt to new tasks with as few as 100 demonstrations. This self-improvement process iteratively finetunes the model and self-generates new data with the finetuned model. Built upon the Gato model [50], RoboCat incorporates the VQ-GAN image encoder [128]. During training, RoboCat predicts not only the next action but also future observations. The effectiveness of the self-improvement process is demonstrated through comprehensive experiments conducted in both simulated and real-world environments under multi-task, multi-embodiment settings.
RT-1 [3], developed by the same team as BC-Z [130], shares similarities with BC-Z but introduces some key distinctions. Notably, RT-1 employs a vision encoder based on the more efficient EfficientNet [132], departing from BC-Z’s use of ResNet18. The language instruction is also encoded using USE [172] and is integrated with image embeddings through FiLM layers [131]. However, RT-1 does not use video as a task instruction, unlike BC-Z. Additionally, RT-1 replaces the MLP action decoder in BC-Z with a Transformer decoder, producing discretized actions. This modification enables RT-1 to attend to past images, enhancing its performance compared to BC-Z.
Q-Transformer [134] extends RT-1 [3] by introducing autoregressive Q-functions. In contrast to RT-1, which learns expert trajectories through imitation learning, Q-Transformer adopts Q-learning methods. Alongside the TD error objective of Q-learning, a conservative regularizer is incorporated to ensure that the maximum value action remains in-distribution. This approach allows Q-Transformer to leverage not only successful demonstrations but also failed trajectories for learning.
III-B3 Multi-modal-instruction Control Policies
VIMA [176] places a significant emphasis on multi-modal prompts and the generalization capabilities of models. By incorporating multi-modal prompts, more specific and intricate tasks can be formulated compared to traditional pure text prompts. VIMA introduces four main types of tasks: object manipulation, visual goal reaching, novel concept grounding, one-shot video imitation, visual constraint satisfaction, visual reasoning. These tasks are often challenging or even infeasible to express using only language prompts. VIMA-Bench has been developed to evaluate across four generalizability levels: placement, combinatorial, novel object, novel task.
MOO [129] extends RT-1 to handle multi-modal prompts. Leveraging the backbone of RT-1, MOO incorporates OWL-ViT [129] to encode images within the prompt. By expanding the RT-1 dataset with new objects and additional prompt images, MOO enhances the generalization capabilities of RT-1. This extension also facilitates new methods of specifying target objects, such as pointing with a finger, and clicking on graphical user interfaces.
III-B4 Control Policies with 3D Vision
Perceiver-Actor (PerAct) [120] represents an advancement in both the observation and action spaces by leveraging 3D voxel representations. This approach offers a robust structural prior for action learning, enabling natural processing of multi-view observations and facilitating data augmentation in 6-DoF. In this framework, the input to the model comprises voxel maps reconstructed from RGBD images, while the output corresponds to the best voxel for guiding the gripper’s movement. By adopting this formulation, PerAct facilitates efficient task learning even with a few demonstrations. On the other hand, Act3D [178] introduces a continuous resolution 3D feature field, with adaptive resolutions based on current task, addressing the computational cost of voxelization. RVT, RVT-2 [122, 123] propose re-rendering images from virtual views and using these images as inputs, rather than directly relying on 3D inputs. RoboPoint [124] leverages VLM to predict affordance points directly on 2D images, which are then projected into 3D space using depth maps.
SceneScript [179] introduces a set of structured language commands that can define an entire scene, specifying the layout and objects. This language-based scene representation distinguishes itself from previous methods—such as meshes, voxel grids, point clouds, or radiance fields—by generating scenes in an autoregressive, token-based fashion. It achieves competitive results against state-of-the-art methods in layout estimation and object detection.
III-B5 Diffusion-based Action Generation
Diffusion-based action generation leverages the success of diffusion models in the field of computer vision [162]. Diffusion policy [161] formulates a robot policy as Denoising Diffusion Probabilistic Models (DDPMs) [162]. This approach incorporates a variety of innovative techniques, including receding horizon control, visual conditioning, and the time-series diffusion transformer. The effectiveness of this diffusion-based visuomotor policy is underscored by its proficiency in multimodal action distributions, high-dimensional action spaces, and training stability.
Scaling Up and Distilling Down (SUDD) [164] presents a framework where LLM guides data generation and subsequently the filtered dataset is distilled into a visuo-linguo-motor policy. This framework achieves language-guided data generation by composing LLM with a suite of primitive robot utilities, such as grasp samplers, motion planners. Then it extends Diffusion Policy [161] through the incorporation of language-based conditioning for multi-task learning. This diffusion-based policy learns from successful trajectories, facilitating the distillation of the filtered dataset.
RDT-1B [171] is a diffusion-based foundation model for bimanual manipulation, also built on DiT. It addresses data scarcity by introducing a unified action format across various robots, enabling pre-training on heterogeneous multi-robot datasets with over 6K trajectories. As a result, RDT scales up to 1.2B parameters and demonstrates zero-shot generalization.
III-B6 Diffusion-based Control Policy with 3D Vision
III-B7 Motion Planner
VoxPoser [154] employs LLM and VLM to create two 3D voxel maps that represent affordance and constraint. It leverages the programming capability of LLM [157], and the perception capability of VLM models: ViLD [155], MDETR [156], OWL-ViT [133], SAM [31]. LLM translates language instructions into executable code, invoking VLM to obtain object coordinates. Based on the composed affordance and constraint maps, VoxPoser employs model predictive control to generate a feasible trajectory for the robot arm’s end-effector. Notably, VoxPoser does not require any training, as it directly connects LLM and VLM for motion planning.
RoboTAP [180] breaks down demonstrations into stages marked by the opening and closing of the gripper. In each stage, RoboTAP uses the TAPIR algorithm to detect active points that track the relevant object from the source pose to the target pose. The path can then be used by classical visual servoing to control the robot. A motion plan is created by stringing together these stages, enabling few-shot visual imitation.
III-B8 LLM-based Control Policies
RT-2 [2] endeavors to harness the capabilities of large multi-modal models in robotics tasks, drawing inspiration from models like PaLI-X [140] and PaLM-E [181]. The approach introduces co-fine-tuning, aiming to fit the model to both Internet-scale visual question answering (VQA) data and robot data. This training scheme enhances the model’s generalizability and brings about emergent capabilities. RT-2 represents an effort to integrate low-level control policy [3] with high-level task planners in pursuit of a more comprehensive robotic system.
RT-X [143] builds upon the previous RT-1 and RT-2 models. These models are re-trained using the newly introduced open-source large dataset, named Open X-Embodiment (OXE), which is orders of magnitude larger than previous datasets. OXE comprises 160,266 tasks from 527 skills with 22 embodiments. The resulting models, RT-1-X and RT-2-X, both outperform their original versions, owing to the more diverse and larger OXE dataset. OpenVLA [144] was later developed as an open-source counterpart to RT-X.
@@ III-B8 LLM-based Control Policies
Most low-level control policies predict actions for the end-effector pose while abstracting away the motion planning module that controls the motion of individual joints with inverse kinematics. While this abstraction facilitates better generalization to different embodiments, it also imposes limitations on dexterity.
@@
$\mathcal{L}_{\text{Cont}}=\sum_{t}\operatorname{MSE}\left(\mathbf{a}_{t},\hat{\mathbf{a}}_{t}\right),$ (1)
where $\operatorname{MSE}(\cdot)$ stands for mean squared error. $\hat{\mathbf{a}}_{t}=\operatorname{argmax}_{a_{t}}\pi_{\theta}(a_{t}|p,s_{\leq t},a_{<t})$ is predicted action while $\mathbf{a}_{t}$ is the action from expert demonstrations.
@@
$\mathcal{L}_{\text{Disc}}=\sum_{t}\operatorname{CE}\left(\mathbf{a}_{t},\hat{\mathbf{a}}_{t}\right),$ (2)
@@
$\begin{split}\mathcal{L}_{\text{{SE}(2)}}&=\operatorname{CE}(\mathbf{a}_{\text{pick}},\hat{\mathbf{a}}_{\text{pick}})+\operatorname{CE}(\mathbf{a}_{\text{place}},\hat{\mathbf{a}}_{\text{place}})\end{split}$ (3)
@@
$\mathcal{L}_{\text{DDPM}}=\operatorname{MSE}\left(\varepsilon^{k},\varepsilon_{\theta}\left(\mathbf{x}^{0}+\varepsilon^{k},k\right)\right)$ (4)
where $\mathbf{x}^{0}$ is the raw example from the dataset and $\varepsilon^{k}$ is a random noise for iteration $k$; $\varepsilon_{\theta}$ is the noise prediction network, i.e., the VLA model.
While discrete action has demonstrated superior performance in RT-1 [3], Octo [165] argues that it leads to early grasping issues. $\textbf{SE}(2)$ action only necessitates the model to predict the two end-effector poses, e.g., pick pose and place pose. While this action type can be predicted in at most two forward passes, it also imposes limitations on action dexterity and generalizability.
LLM vs non-LLM
While LLM-based control policies can greatly enhance instruction-following abilities because LLM can better parse user intentions, concerns arise regarding their training cost and deployment speed. Slow inference speed, in particular, can significantly impact performance in dynamic environments, as changes in the environment may occur during LLM inference.
@@
RT-1 [3] inspired a series of “Robotic Transformer” models. Preceding RT-1 was BC-Z [130], which solely utilized MLP layers for action prediction. Subsequent to RT-1, several works emerged, each introducing new capabilities. MOO [129] adapted RT-1 to accommodate multi-modal prompts. RT-Trajectory [135] enabled RT-1 to process trajectory sketches as prompts. Q-Transformer [134] utilized Q-learning to train RT-1. RT-2 [2], based on ViT and LLM, introduced a completely different architecture from RT-1. RT-X [143] retrained RT-1 and RT-2 with a significantly larger dataset, resulting in improved performance. The Transformer backbone surpasses previous RNN backbones [110, 111, 112] by harnessing the higher capacity of Transformers to absorb larger robot datasets. Based on RT-2, RT-H [142] introduces action hierarchies for better data sharing.

### v4

@@ III-B1 Non-Transformer Control Policies
BC-Z [113] processes two types of task instructions: a language instruction or a human demonstration video. The environment is presented to the model in the form of an RGB image. Then the instruction embedding and the image embedding are combined through the FiLM layer [148], culminating in the generation of actions. This conditional policy is asserted to exhibit zero-shot task generalization to unseen tasks.
@@
UniPi [128] treats the decision-making problem as a text-conditioned video generation problem. To predict actions, UniPi generates a video based on a given text instruction, and actions are extracted from the frames of the video through inverse dynamics. This innovative policy-as-video formulation offers several advantages, including enhanced generalization across diverse robot tasks and the potential for knowledge transfer from internet videos to real robots.
III-B2 Transformer-based Control Policies
Since the introduction of Transformers, control policies have converged to similar Transformer-based architectures.
RoboCat [109] proposes a self-improvement process designed to enable an agent to rapidly adapt to new tasks with as few as 100 demonstrations. This self-improvement process iteratively finetunes the model and self-generates new data with the finetuned model. Built upon the Gato model, RoboCat incorporates the VQ-GAN image encoder [110]. During training, RoboCat predicts not only the next action but also future observations. The effectiveness of the self-improvement process is demonstrated through comprehensive experiments conducted in both simulated and real-world environments under multi-task, multi-embodiment settings.
RT-1 [114], developed by the same team as BC-Z, shares similarities with BC-Z but introduces some key distinctions. Notably, RT-1 employs a vision encoder based on the more efficient EfficientNet [115], departing from BC-Z’s use of ResNet. However, RT-1 does not use video as a task instruction. Additionally, RT-1 replaces the MLP action decoder in BC-Z with a Transformer decoder, producing discretized actions. This modification enables RT-1 to attend to past images, enhancing its performance compared to BC-Z.
Q-Transformer [117] extends RT-1 by introducing autoregressive Q-functions. In contrast to RT-1, which learns expert trajectories through imitation learning, Q-Transformer adopts Q-learning methods. Alongside the TD error objective of Q-learning, a conservative regularizer is incorporated to ensure that the maximum value action remains in-distribution. This approach allows Q-Transformer to leverage not only successful demonstrations but also failed trajectories for learning.
III-B3 Control Policies for Multimodal Instructions
Multimodal instruction enables new ways to specify tasks, such as through demonstrations, by naming novel objects, or by pointing with a finger.
VIMA [150] places a significant emphasis on multimodal prompts and the generalization capabilities of models. By incorporating multimodal prompts, more specific and intricate tasks can be formulated compared to traditional pure text prompts. VIMA introduces four main types of tasks: object manipulation, visual goal reaching, novel concept grounding, one-shot video imitation, visual constraint satisfaction, visual reasoning. These tasks are often challenging or even infeasible to express using only language prompts. VIMA-Bench has been developed to evaluate across four generalizability levels: placement, combinatorial, novel object, novel task.
MOO [111] extends RT-1 to handle multimodal prompts. Leveraging the backbone of RT-1, MOO incorporates OWL-ViT to encode images within the prompt. By expanding the RT-1 dataset with new objects and additional prompt images, MOO enhances the generalization capabilities of RT-1. This extension also facilitates new methods of specifying target objects, such as pointing with a finger, and clicking on graphical user interfaces.
III-B4 Control Policies with 3D Vision
PerAct [103] represents an advancement in both the observation and action spaces by leveraging 3D voxel representations. This approach offers a robust structural prior for action learning, enabling natural processing of multi-view observations and facilitating data augmentation in 6-DoF. In this framework, the input to the model comprises voxel maps reconstructed from RGBD images, while the output corresponds to the best voxel for guiding the gripper’s movement. By adopting this formulation, PerAct facilitates efficient task learning even with a few demonstrations. On the other hand, Act3D [104] introduces a continuous resolution 3D feature field, with adaptive resolutions based on current task, addressing the computational cost of voxelization. RVT, RVT-2 [105, 106] propose re-rendering images from virtual views of the scene point cloud and using these images as inputs, rather than directly relying on 3D inputs.
III-B5 Diffusion-based Control Policies
Diffusion-based action generation leverages the success of diffusion models in the field of computer vision.
Diffusion Policy [130] formulates a robot policy as a DDPM [131]. This approach incorporates a variety of techniques, including receding horizon control, visual conditioning, and the time-series diffusion transformer. The effectiveness of this diffusion-based visuomotor policy is underscored by its proficiency in multimodal action distributions, high-dimensional action spaces, and training stability.
SUDD [133] presents a framework where an LLM guides data generation, and subsequently, the filtered dataset is distilled into a visuo-linguo-motor policy. This framework achieves language-guided data generation by composing an LLM with a suite of primitive robot utilities, such as grasp samplers and motion planners. It then extends Diffusion Policy by incorporating language-based conditioning for multi-task learning and facilitates the distillation of the filtered dataset.
RDT-1B [139] is a diffusion-based foundation model for bimanual manipulation, also built on DiT. It addresses data scarcity by introducing a unified action format across various robots, enabling pretraining on heterogeneous multi-robot datasets with over 6K trajectories. As a result, RDT scales up to 1.2B parameters and demonstrates zero-shot generalization.
III-B6 Diffusion-based Control Policies with 3D Vision
III-B7 Control Policies for Motion Planning
Motion planning involves decomposing movement tasks into discrete waypoints while satisfying constraints such as obstacle avoidance and kinematic limits.
VoxPoser [125] employs LLM and VLM to create two 3D voxel maps that represent affordance and constraint. It leverages the programming capability of LLMs, and the perception capability of VLMs. LLM translates language instructions into executable code, invoking VLM to obtain object coordinates. Based on the composed affordance and constraint maps, VoxPoser employs model predictive control to generate a feasible trajectory for the robot arm’s end-effector. Notably, VoxPoser does not require any training, as it directly connects LLM and VLM for motion planning.
RoboTAP [152] breaks down demonstrations into stages marked by the opening and closing of the gripper. In each stage, RoboTAP uses the TAPIR algorithm to detect active points that track the relevant object from the source to the target pose. The path can then be used by visual servoing to control the robot. A motion plan is created by stringing together these stages, enabling few-shot visual imitation.
III-B8 Control Policies with Point-based Action
Recent research has explored leveraging the capabilities of VLMs to select or predict point-based actions—a cost-effective alternative to building full VLAs.
PIVOT [153] casts robotics tasks as visual question answering, leveraging VLMs to select the best robot action from a set of visual proposals. Visual proposals are annotated in the form of keypoints on images. The VLM is iteratively prompted to refine them until the best option is identified.
RoboPoint [107] finetunes VLM using the task of spatial affordance prediction, which is to point at where to act on the image. These affordance points on 2D images are subsequently projected into 3D space using depth maps, forming the predicted robot action.
ReKep [51] is a constraint function that maps 3D keypoints in the scene to a numerical cost. Robotics manipulation tasks can be represented as a sequence of ReKep constraints, which are produced with large vision models and VLMs. Consequently, robot actions can be obtained by solving the constrained optimization problem.
III-B9 Large VLA
Large VLA is equivalent to the original VLA definition proposed by RT-2 [2], as illustrated in Figure 2a. This terminology is analogous to the distinction between LLMs and general language models, or between large VLMs and general VLMs.
RT-2 [2] endeavors to harness the capabilities of large multimodal models in robotics tasks, drawing inspiration from models like PaLI-X and PaLM-E. The approach introduces co-fine-tuning, aiming to fit the model to both Internet-scale visual question answering (VQA) data and robot data. This training scheme enhances the model’s generalizability and brings about emergent capabilities.
RT-X [143] builds upon the previous RT-1 and RT-2 models. These models are re-trained using the newly introduced open-source large dataset, named Open X-Embodiment (OXE), which is orders of magnitude larger than previous datasets. The resulting models, RT-1-X and RT-2-X, both outperform their original versions.
OpenVLA [50] was later developed as an open-source counterpart to RT-2-X. They additionally explored efficient fine-tuning methods, including LoRA and model quantization.
$\pi_{0}$ [147] proposes a flow-matching architecture for transforming VLMs into VLAs. By incorporating an additional action expert based on the mixture-of-experts framework, it effectively inherits the Internet-scale knowledge in the VLM while extending its capabilities to address robotic tasks.
@@ III-B9 Large VLA
Figure 5: Three common vision-language fusion methods of Transformer-based control policies. FiLM layers (Hadamard product $\odot$) are used in RT-1 models. Some utilize cross-attention to condition on instructions. Concatenation ($\oplus$) is the prevailing method in LVLAs.
@@
Most low-level control policies predict actions for the end-effector pose while abstracting away the motion planning module that produces more fine-grained motions. While this abstraction facilitates better generalization to different embodiments, it also imposes limitations on dexterity.
@@
$\mathcal{L}_{\text{Cont}}=\sum_{t}\operatorname{MSE}\left(a_{t},\hat{a}_{t}\right),$ (4)
where $\operatorname{MSE}(\cdot)$ stands for mean squared error. $a_{t}$ is the action annotation from expert demonstrations.
@@
$\mathcal{L}_{\text{Disc}}=\sum_{t}\operatorname{CE}\left(a_{t},\hat{a}_{t}\right),$ (5)
@@
$\begin{split}\mathcal{L}_{SE(2)}&=\operatorname{CE}(a_{\text{pick}},\hat{a}_{\text{pick}})+\operatorname{CE}(a_{\text{place}},\hat{a}_{\text{place}})\end{split}$ (6)
@@
$\mathcal{L}_{\text{DDPM}}=\operatorname{MSE}\left(\varepsilon^{k},\varepsilon_{\theta}(a_{t}+\varepsilon^{k},k)\right)$ (7)
where $\varepsilon^{k}$ is a random noise for iteration $k$; $\varepsilon_{\theta}$ is the noise prediction network, i.e., the VLA model.
While discrete action has demonstrated superior performance in RT-1 [114], Octo [134] argues that it leads to early grasping issues. $SE(2)$ actions require the model to predict only the pick and place end-effector poses, which is sufficient for many tabletop manipulation tasks. However, more complex tasks—such as “pouring water into a cup”—may require additional DoFs, thereby necessitating $SE(3)$ actions [51]. Although point-based actions can be coarse-grained, they are more easily obtained from VLMs in a zero-shot manner [51].
Table IV: High-level task planners. VL: vision-language fusion method.
Model V-Encoder Lang Model VL Control Policy Train Robotic Tasks
SayCan [10] - PaLM, FLAN - BC-Z, MT-Opt No Real-Navi&Mani (Everyday): office kitchen
[…]
(SL)3 [155] - T5-small - seq2seq Yes Sim-Navi: ALFRED
EmbodiedGPT [156] EVA ViT-G/14 [157] LLaMA-7B Embodied-former MLP Yes Sim-Mani: Franka Kitchen, Meta-World
PaLM-E [11] ViT, OSRT [158] PaLM Concat Interactive Language, RT-1 Yes Sim: TAMP; Sim&Real-Mani: Lang.-Table; Real: SayCan setup
LEO [159] ConvNeXt [160], PointNet++ [161] Vicuna-7B Concat MLP Yes Sim-Mani: CLIPort task subset; Sim-Navi: Habitat-web
3D-LLM [57] CLIP, 3D-CLR [162], ConceptFusion [163] Flamingo, BLIP-2 Concat DD-PPO [164] Yes Sim-Navi: Habitat
ShapeLLM [165] ReCon++ LLaMA Concat - Yes 3D MM-Vet Benchmark
Inner Monologue [9] MDETR, CLIP InstructGPT, PaLM Language CLIPort, MDETR, SayCan policies No Sim&Real-Mani: Tabletop Rearrangement; Real-Navi&Mani: Kitchen Mobile Mani.
LLM-Planner [166] HLSM GPT-3 Language HLSM Yes Sim-Navi: ALFRED
LID [167] CNN GPT-2 Language & Embed - Yes Sim: VirtualHome
SMs [168] CLIP, ViLD GPT-3, RoBERTa [169] Language & Code CLIPort No Sim-Mani: pick, place
ConceptGraphs [170] SAM, CLIP, LLaVA [8] GPT-4 Code (JSON) API No Sim: AI2-THOR; Real (Spot Arm): pick, place
ProgPrompt [171] ViLD GPT-3 Code API No Sim: VirtualHome
ChatGPT for Robotics [172] YOLOv8 [173] ChatGPT Code API No Real-Navi: drone flight; Sim-Navi: AirSim, Habitat
CaP [174] ViLD, MDETR GPT-3, Codex Code API No Sim-Mani & Sim-Navi: pick, place, etc.
DEPS [175] CLIP ChatGPT Code MC-Controller, Steve-1 No Sim: Minecraft
@@
RT-1 [114] inspired a series of “Robotic Transformer” models. The Transformer backbone surpasses previous RNN backbones by harnessing the higher capacity of Transformers to absorb larger robot datasets. Preceding RT-1 was BC-Z, which solely utilized MLP layers for action prediction. Subsequent to RT-1, several works emerged, each introducing new capabilities. MOO adapted RT-1 to accommodate multimodal prompts. RT-Trajectory enabled RT-1 to process trajectory sketches as prompts. Q-Transformer utilized Q-learning to train RT-1. RT-2, based on ViT and LLM, introduced a completely different architecture from RT-1. RT-X retrained RT-1 and RT-2 with a significantly larger dataset, resulting in improved performance. Based on RT-2, RT-H [142] introduces action hierarchies for better data sharing.
LVLA vs VLA
While LVLAs can greatly enhance instruction-following abilities because they can better parse user intentions, concerns arise regarding their training cost and deployment speed. Slow inference speed, in particular, can significantly impact performance in dynamic environments, as changes in the environment may occur during inference. Therefore, several methods have been proposed to improve efficiency. TinyVLA [176] focused on inference speed and data efficiency through smaller VLM and diffusion head for robot action. DeeR-VLA [177] proposes only partially activating the model with dynamic inference with early-exit.
IV Task Planners
A high-level task planner $\pi_{\phi}$ aims to divide a complex task $\ell$ into a sequence of subtasks $p_{i}$ (i.e., task plan), each serving as a language instruction to low-level control policies $\pi_{\theta}$:
$\begin{split}\left[p_{1},p_{2},\dots,p_{N}\right]\sim\pi_{\phi}(\ell,s_{t}).\end{split}$ (8)
This process is often referred to as task decomposition and is also related to the task and motion planning (TAMP) setting. Ideally, task plans should also incorporate optimal scheduling for these subtasks. When equipped with task planners, VLAs can complete more complex, long-horizon tasks, as illustrated in Figure 4. Table IV outlines important details about task planners.
IV-A Monolithic Task Planners
A single LLM or multimodal LLM (MLLM) can typically generate task plans by employing a tailored framework or through finetuning on embodied datasets. We refer to these as monolithic models.
IV-A1 Grounded Task Planners
Grounded task planning involves generating high-level actions while considering whether they can be executed by low-level control policies.
[…]
(SL)3 [155] is a learning algorithm that alternates between three steps: segmentation, labeling, and parameter update. In the segmentation step, high-level subtasks are aligned with low-level actions, subtask descriptions are then inferred in the labeling step, and finally, the network parameters are updated. This approach enables a hierarchical policy to discover reusable skills with sparse natural language annotations.
IV-A2 End-to-end Task Planners
Similar to LVLAs, task planners can be implemented as end-to-end multimodal LLMs, leveraging their Internet-scale knowledge for task planning. Clearly, they are also monolithic models.
PaLM-E [11] integrates ViT and PaLM to create a large embodied multimodal language model capable of performing high-level embodied reasoning tasks. Based on perceived images and high-level language instructions, PaLM-E generates a text plan that serves as instructions for low-level robotic policies. In a mobile manipulation environment, they map the generated plan to executable low-level instructions with SayCan [10]. As the low-level policy executes actions, PaLM-E can also replan based on changes in the environment. With PaLM as its backbone, PaLM-E can handle both normal visual question answering (VQA) tasks, along with the additional embodied VQA tasks.
IV-A3 End-to-end Task Planners with 3D Vision
Some task planners also explore the use of 3D vision. Because the majority of current MLLMs deal with images as visual inputs, they require architectural changes to incorporate 3D visual inputs, and therefore they are usually end-to-end models.
LEO [159] uses a two-stage training strategy to integrate a point cloud encoder with an LLM: the first stage focuses on 3D vision-language alignment, followed by the second stage, which involves 3D vision-language-action instruction tuning. LEO demonstrates proficiency not only in 3D question-answering tasks but also in manipulation, navigation, and task planning.